OpenStudy (anonymous):

What is the graph of the equation? –5x + y = –3 (1 point)

OpenStudy (anonymous):

Well, to find the graph you have to put it into slope intercept form, which is y=mx+b

OpenStudy (anonymous):

so... it's y=5x-3

OpenStudy (anonymous):

And then to find the x point, substitute y with 0

OpenStudy (anonymous):

0=5x-3 +3 +3 --------- 3=5x /5 /5 --------- 3/5=x

OpenStudy (anonymous):

And do the same for y

OpenStudy (anonymous):

y=3

OpenStudy (anonymous):

So your points are (3/5, 3)

OpenStudy (anonymous):

Do you understand it? :3 Most important is that you get it.

OpenStudy (anonymous):

yup